my email address is shamimkhan.khan52@gmail.com
and i want to know about a question who is given below:-
1+3+5+7+11+13+15+17+19=100
find 5 words whose sum is 50 But take 5 words from above line

Answer by Theo(13342)

you are missing a 9 in your sequence.
the sequence should be:
1 + 3 + 5 + 7 + 9 + 11 + 13 + 15 + 17 + 19 = 100
that's 10 numbers in total.
you are looking for 5 numbers in total to be equal to 1/2 of 100.
100 is an even number.
50 is an even number.
if you add 2 odd numbers together you will always get an even number.
that's why you can add 10 odd numbers together to get an even number.
when you reduce that by half, you are looking for 5 odd numbers that add up to an even number.
that can't be done.
adding an odd number of odd numbers will always result in an odd number.
you will not be able to get 5 numbers in this sequence to total up to 50 because 50 is an even number and you are wanting to add an odd number of odd numbers in order to get it.
consider this.
take your 10 numbers and pair them with each other.
(1+3) = 4
(5 + 7) = 12
(9 + 11) = 20
(13 + 15) = 28
(17 + 19) = 36
now take any of 5 of these numbers and pair them together.
(5+7) = 12
(13 + 15) = 28
the next number is all by itself and is odd.
you will never get 50 out of any 5 of these numbers.